When i was 4 years, my sister was half my age. now my sister is 30 years old. how old am i now?
Answer by rajagopalan(174)

I was 4 years,
my sister was half my age.
at that point my sister's age =4x0.5=2
Age difference is 4-2=2
now my sister is 30 years old.
My age = sis age+2=30+2=32.
Now My age = 32.
Bravo..